Sales Mix and Break-Even Analysis

Michael Company has fixed costs of $496,640. The unit selling price, variable cost per unit, and contribution margin per unit for the company’s two products follow:

Product Model Selling Price Variable Cost per Unit Contribution Margin per Unit
Yankee $180 $80 $100
Zoro 140 120 20

The sales mix for products Yankee and Zoro is 55% and 45%, respectively. Determine the break-even point in units of Yankee and Zoro.

a. Product Model Yankee  units

b. Product Model Zoro  units
Answer #1

Weighted average contribution margin per unit

= (100*55%) + (20*45%) = 64

Overall break even point = Fixed cost/weighted average contribution margin

= 496,640/64 = 7760 units

Product model Yankee units = (7760*55%) 4268
Product model Zoro units = (7760*45%) 3492